OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = 3x − 4, find f[g(3)]. 1. 11 2. 17 3. 5 4. 7

OpenStudy (anonymous):

\[g(3)=3\times 3-4=9-4=5\] \[f(g(3))=f(5)=2\times 5+1=10+1=11\]

OpenStudy (anonymous):

yes what he said